Given the arithmetic sequence 2, 6, ...
If t(0) is the first term find a formula for t(n).

Answer by psbhowmick(878)

t(0) = 2
Common Difference: d = 6-2=4
The n-th term: t(n) = t(0) + (n-1)d = 2+(n-1)4 = 4n-2